OpenStudy (anonymous):

find the distance between the parallel lines with the given equations. 2x-3y+4=0 and y=2/3x+5

OpenStudy (anonymous):

you might have some formula for this right?

OpenStudy (anonymous):

first equation is 2x - 3y = -4 and second one is 3y = 2x + 15 => 2x - 3y = -15 { we just took those equation in a*x + b*y = constant.

OpenStudy (anonymous):

now we will ensure that the coefficients of x and y in both equations are same { this is important }, if they are not equal, you would've to multiply either of equations with some number. here they are already there, it's ok

OpenStudy (anonymous):

then in this form the difference between the two constants -4 and -15 is the distance between them. here -4 - ( -15) = -4 + 15 = 11 units.

OpenStudy (academicgurusinc):

y= (2/3)(2) + (4/3) =8/3 y= (2/3)(2) +5 = 19/3 19/3 - 8/3 =8/3

OpenStudy (academicgurusinc):

Because the two lines are parallel you need only find the y values of each of the two equations and then subtract the two.

OpenStudy (academicgurusinc):

We hope this helps. Regards, Academic GurusInc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)

OpenStudy (anonymous):

thank you so much guys :)

OpenStudy (anonymous):

now we got 11 and the coefficient of y is 3,so we need to divide by 3. thus the answer 11/3.

OpenStudy (anonymous):

the answer is 11sqrt13/13 because were working with the "normal"

OpenStudy (anonymous):

@AcademicGurusInc your explanation is smooth +1, but you oversaw a little error, 19/3 - 8/3 is 11/3 not 8/3 :P

OpenStudy (anonymous):

as for the second question i have two more lines 4x-y+1=0 and 4x-y-8=0

OpenStudy (academicgurusinc):

@SandeepReddy thank you for the correction. @Sophie_dofi sorry for the oversight. I hope you were able to understand the concept behind the solution.

OpenStudy (academicgurusinc):

AS for the second question: The idea is very similar. First decide on the relationship between the two lines. Are they also parallel? (Hint: take a look at the slopes)

OpenStudy (anonymous):

yea i understood it :) i should have mentioned I'm working with normal it was my bad

OpenStudy (anonymous):

yea they are parellel

OpenStudy (academicgurusinc):

Ok no worries. So are you okay to solve the second question or do you need my help?

OpenStudy (anonymous):

well im used to finding the normal with just one equation im a bit lost with two

OpenStudy (academicgurusinc):

Ok, so let's take it step by step. You found that the two lines are parallel because they both have the same slope (4). That means regardless which x value you choose, you can find the distance by solving for the respective y values. Try using the x value 3, for example. What y values do you get for the each of the two equations?

OpenStudy (anonymous):

-y=-1-4x? -y=8-4x sorry i dont have paper so thats probably wrong actually im pretty sure i didnt do that correctly

OpenStudy (academicgurusinc):

You're on the right track. You now need to multiply the equation through by -1 to remove the negative sign. So they the two equations should look like this: 4x-y+1=0 -----> y =4x + 1 4x-y-8=0 ------> y = 4x-8

OpenStudy (academicgurusinc):

now solve for y using x=3

OpenStudy (anonymous):

y= 13 y=4

OpenStudy (academicgurusinc):

perfect. now subtract the two. What do you find

OpenStudy (anonymous):

you mean subtract them from each other? 9

OpenStudy (academicgurusinc):

That is correct. The distance is 9

OpenStudy (anonymous):

so 9 squt17/17?

OpenStudy (academicgurusinc):

so 9 squt17/17? Sorry, are you trying to find the distance of the two lines you mentioned? I'm not sure of where you got your result from

OpenStudy (academicgurusinc):

If you are trying to find the distance between the two lines, 9 is correct.

OpenStudy (anonymous):

i used normal form

OpenStudy (academicgurusinc):

Unfortunately I will have to leave. If there the question was asking for something other than the distance please feel free to send us a message. We hope you found the solution helpful. Regards, Academic Gurus Inc.

OpenStudy (anonymous):

thanks :)

OpenStudy (freckles):

Here is the idea: Say we have the two parallel lines: y=mx+b y=mx+a We need to find a perpendicular line to both. The slope of a perpendicular line will be -1/m. So the easier perpendicular line we can choose is y=-1/m x . Now we just need to find the solutions to the two systems: y=mx+b y=-1/m x and also y=mx+a y=-1/m x So let's solve the first system by substitution: y=mx+b y=-1/m x so we have -1/m x=mx+b Solving for x.... Subtract mx on both sides \[- \frac{1}{m}x- mx =b \\ -\frac{1}{m}x-\frac{m^2}{m}x=b\\ \frac{-1-m^2}{m}x=b \\ x=\frac{mb }{-1-m^2} \\ x=\frac{-mb}{1+m^2} \] \[y=\frac{-1}{m} \cdot \frac{-mb}{1+m^2}=\frac{b}{1+m^2}\] We can solve the other system in a similar way: y=mx+a y=-1/m x \[x=\frac{-ma}{1+m^2}\] \[y=\frac{a}{1+m^2}\]  Now find the distance between these two points: \[d=\sqrt{(\frac{-mb}{1+m^2}-\frac{-ma}{1+m^2})^2+(\frac{b}{1+m^2}-\frac{a}{1+m^2})^2} \\ d=\sqrt{(\frac{-mb+ma}{1+m^2})^2+(\frac{b-a}{1+m^2})^2} \\ d=\sqrt{(\frac{m(-b+a)}{1+m^2})^2+(\frac{b-a}{1+m^2})^2} \\ d=\sqrt{ m^2 (\frac{(-1)(b-a)}{1+m^2})^2+(\frac{b-a}{1+m^2})^2} \\ d=\sqrt{m^2(-1)^2 (\frac{b-a}{1+m^2})^2+(\frac{b-a}{1+m^2})^2}\] \[d=\sqrt{m^2(\frac{b-a}{1+m^2})^2+(\frac{b-a}{1+m^2})^2} \\ d=\sqrt{(\frac{b-a}{1+m^2})^2(m^2+1)} \\ d=\sqrt{(\frac{b-a}{1+m^2})^2 } \sqrt{m^2+1} \\ d=|\frac{b-a}{1+m^2}| \sqrt{m^2+1}\] d is the shortest distance between the two parallel lines y=mx+a and y=mx+b